OpenStudy (anonymous):

determine the midpoint of the line segment with each set of given endpoints. (8,0) and (4,6)

OpenStudy (madhu.mukherjee.946):

mid pt =y2-y1/x2-x1 =6-0/4-8=6/-4=-1.5